OpenStudy (anonymous):

Let f() = 3x+cosx Evaluate f^-1(1)

OpenStudy (anonymous):

\[ 1=3x+\cos x \]

OpenStudy (anonymous):

How do you show it's one - to - one? Graphing ?

OpenStudy (anonymous):

What is one to one?

OpenStudy (anonymous):

Can't remember it entirely but it has something to with the the range I believe

OpenStudy (anonymous):

hits the graph at most once

OpenStudy (anonymous):

No, I know what one to one means. Are you trying to prove that \(f(x)\) is one to one?

OpenStudy (anonymous):

ah yes lol

OpenStudy (anonymous):

The derivative is \[ f'(x) = 3-\sin(x) \]Since \(\min\{\sin(x)\}=-1\) then\[ \min\{f'(x)\} = 3-1=2 \]

OpenStudy (anonymous):

oooooh

OpenStudy (anonymous):

\[ \forall x\quad f'(x)>2\implies f(x) \text{ monotonically increasing} \]

OpenStudy (anonymous):

Suppose that it was not one to one. Then there must exists two distinct numbers \(a,b\) such that \(f(a)=f(b)\). But if \(a<b\) then \(f(a)<f(b)\) and if \(a>b\) then \(f(a)>f(b)\) which is a contradition.

OpenStudy (anonymous):

So how do you actually solve this inverse? I don't think x can be solved in terms of y?

OpenStudy (anonymous):

got ya

OpenStudy (anonymous):

It doesn't have to necessarily be an elementary function I suppose.

OpenStudy (anonymous):

So you're basically using inspection then lol

OpenStudy (anonymous):

Well finding roots was never a trivial thing in mathematics.

OpenStudy (anonymous):

You need to find the roots of \[ 3x+\cos x - 1=0 \]

OpenStudy (anonymous):

f^-1(1) = x means f(x) = 1 3x+cosx = 1 f(0)= 3(0)+cos(0) = 1

OpenStudy (anonymous):

Yes, always try easy angles.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

f^-1(1) therefore = 0

OpenStudy (anonymous):

Also take node of the amplitude of the trig function. That narrows things down a lot.

OpenStudy (anonymous):

For example, we knew that \(x<1\) because that \(3x\) was very large.

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

If it wasn't 0, it was going to be a very nasty number because trig functions need transcendental numbers just to get algebraic outputs.

OpenStudy (anonymous):

Thanks izzy :D

OpenStudy (anonymous):

I mean wio ;)